# Erasing Self-Supervised Learning Backdoor by Cluster Activation Masking

### v1

@@ C.2 More Trigger Attention Maps of Different Target Categories
In this section, we show more trigger attention maps of different target categories to further investigate the search processes of two defenders, PoisonCAM and PatchSearch. Both methods generate trigger attention maps for images and select the hottest regions as the candidate triggers. PatchSearch adopts Grad-CAM [45] to compute the trigger attention, while we propose a Cluster Activation Masking method to solve this problem. In Figures 8, 9, 10, we visualize more cases of trigger attention maps generated by PoisonCAM and PatchSearch on ImageNet-100 (poison rate $0.5\%$) with the target categories “tabby cat”, “ambulance” and “pickup truck”. Our method can always accurately locate the backdoor triggers contained in these images, while PatchSearch focuses on more dispersed and maybe irrelevant regions. These results further demonstrate that our PoisionCAM can significantly improve the detection accuracy of the injected backdoor trigger, which is fundamental for defending against SSL backdoor attacks.
C.3 Analysis of Backdoored SSL Process
In our assessment of the backdoored SSL process on ImageNet-100 with a poison rate of $0.5\%$, we conduct a thorough investigation into the correlation between the overall model performance, as measured by Clean Data Acc, and the attack effectiveness, as measured by the Patched Data FP. As shown in Table 3, our findings indicate that in the initial epochs ($epoch\leq 40$), attack effectiveness has a strong correlation with overall model performance. Although this correlation is not as stringent in later epochs, the attack effectiveness still maintains its overall magnitude. These observations suggest that, when training a model for defense, achieving convergence may not be necessary. Instead, it is sufficient to train the model until the attack effectiveness is comparable to that of the fully trained model in overall magnitude.
C.4 Hyper-parameter Sensitivity Study
Refer to caption (a) IoU
Refer to caption (b) CR
Refer to caption (c) Recall
Refer to caption (d) Precision
@@ C.4 Hyper-parameter Sensitivity Study
Refer to caption (a) Acc
Refer to caption (b) IoU
Refer to caption (c) CR
Refer to caption (d) Time
Figure 12: Hyper-parameter sensitivity study on the cluster count of $l$ on ImageNet-100 (poison rate 0.5%, target category “rottweiler”). (a-c) Acc, IoU, and CR of the detected candidate triggers. (d) Total processing time and average processing time per image of candidate trigger detection.
In this section, we analyze hyper-parameter sensitivity studies on window size $w$ of masks and the cluster count of $l$ on ImageNet-100 with poison rate 0.5%, target category “rottweiler”.
@@
Number $l$ of clusters. As shown in Figure 12, we vary the cluster number $l$ in $k$-means algorithm from 100 to 2000 and have the following observations: (1) While increasing $l$, the Acc of the detected candidate triggers first increases and then becomes relatively stable, and the IoU and CR increase first and then fluctuate. These results demonstrate that a large cluster count of $l$ can detect more candidate images with poisonous triggers. (2) The total processing time consistently increases and the average processing time performs relatively stable while the cluster count of $l$ increases. The system will spend more calculating time with a larger cluster count of $l$. To conclude, an appropriate value of $l$ should be chosen to balance the performance and time costs. We empirically set $l=1000$.
C.5 Analysis of Failure Examples
To conduct a comprehensive analysis of our proposed method and promote further advances in related areas, we analyze the typical failure examples in this section.
C.5.1 Undetected Poisoned Samples
Refer to caption Oversized Trigger
Refer to caption Undersized Trigger
Figure 13: The Poisoned Classifier Misses Examples during the Filtering Stage.
Figure 14: During the process of searching for candidate triggers, there are a few instances of search errors, which can be categorized into three types. (a) There are cases where some pixels in the image closely resemble the actual trigger. (b) The trigger in the image occupies a significant portion. (c) Some inherent features present in the image itself can have a significant impact on the model, acting as a natural trigger, such as a pattern resembling a beehive.
We explore the common characteristics among these undetected poisoned samples with the poison classifier discussed in Section 4.4. As shown in Figure 13, the poison classifier wrongly recognizes these oversized and undersized poisonous images as clean images. The results show the limitation of the poison classifier.
C.5.2 Trigger Search Errors
Though our PoisonCAM achieves high recall and precision in the poisonous image search stage discussed in Sections 4.2 and 4.3, there still are a few search errors. As shown in Figure 14, we can have the following observations: (1) Image resembling the trigger. When some pixels in the candidate image are similar to the backdoor trigger. (2) Image corrupted by the trigger. Some triggers are oversized when pasted to form poisonous images, which are hard to search. (3) Image with the analogous trigger. The features of some images like honeycombs are similar to the backdoor trigger.

### v2

@@ VI-E Analysis of Trigger Attention Maps
To further investigate the process of two defenders, PoisonCAM and PatchSearch, we conduct a qualitative analysis of the generated trigger attention maps. Both methods generate trigger attention maps for images and select the hottest regions as the candidate triggers. PatchSearch adopts Grad-CAM [18] to compute the trigger attention, while we propose a Cluster Activation Masking method in Section IV-B to solve this problem. Therefore, we visualize the cases of trigger attention maps generated by PoisonCAM and PatchSearch on ImageNet-100 (poison rate 0.5%) with the target category “rottweiler” in Figure 6. Our method can always accurately locate the backdoor triggers contained in these images, while PatchSearch focuses on more dispersed and maybe irrelevant regions. In Figures 7, 8, 9, we visualize more cases of trigger attention maps generated by PoisonCAM and PatchSearch on ImageNet-100 (poison rate $0.5\%$) with the target categories “tabby cat”, “ambulance” and “pickup truck”. Our method can always accurately locate the backdoor triggers contained in these images, while PatchSearch focuses on more dispersed and maybe irrelevant regions. These results further demonstrate that our PoisionCAM can significantly improve the detection accuracy of the injected backdoor trigger, which is fundamental for defending against SSL backdoor attacks.
VI-F Defend with Data Augmentation
Except for cleaning poisoned data, data augmentations can also relieve backdoor attacks since they may break triggers in training, Moreover, these two types of methods can be integrated in defending. Therefore, we conducted defense experiments with complex i-CutMix [62] augmentation on datasets containing 0.5% and 1.0% poisoned samples. We experiment with i-CutMix, Patch Search, our PoisonCAM, and their combinations.
The experimental results are shown in Figure 10, from which we have the following observations:
  1. 1.
    i-CutMix outperforms PatchSearch in ACC and can achieve sufficiently low ASR. This shows that appropriate data augmentation can relieve backdoor attacks on SSL models.
  2. 2.
    Combined with i-CutMix, the performance of both PatchSearch and PoisonCAM can be improved, which shows the complementarity of two defense manners.
  3. 3.
    Our PoisonCAM with i-CutMix consistently achieves the highest ACC with sufficiently low ASR on two datasets, which further verifies the effectiveness of our method in practice.
(a) IoU
(b) CR
(c) Recall
(d) Precision
@@ VI-F Defend with Data Augmentation
(a) Acc
(b) IoU
(c) CR
(d) Time
Figure 13: Hyper-parameter sensitivity study on the cluster count of $l$ on ImageNet-100 (poison rate 0.5%, target category “rottweiler”). (a-c) Acc, IoU, and CR of the detected candidate triggers. (d) Total processing time and average processing time per image in candidate trigger detection phase.
VI-G Hyper-parameter Sensitivity Study
In this section, we analyze the hyper-parameter sensitivity on ImageNet-100 (poison rate 0.5%) with the target category “rottweiler”.
Number $B$ of masks. As demonstrated in Figure 11, we have the following observations: (1) While increasing the number $B$ of masks, IoU, CR of the detected candidate triggers and F1-score of the trained poison classifier first increase and then become relatively stable, which shows that our method can effectively locate the triggers with appropriate large $B$. The searching time increases with more masks to calculate. (2) The average processing time consistently increases while the number $B$ of masks increases. Considering the worst situation where all masks are processed serially, the slope of the time curve will tend towards a constant. To conclude, an appropriate value of $B$ should be chosen to balance the performance and time costs. We empirically set $B=256$ in our method.
@@ VI-G Hyper-parameter Sensitivity Study
Number $l$ of clusters. As shown in Figure 13, we vary the cluster number $l$ in $k$-means algorithm from 100 to 2000 and have the following observations: (1) While increasing $l$, the Acc of the detected candidate triggers first increases and then becomes relatively stable, and the IoU and CR increase first and then fluctuate. These results demonstrate that a large cluster count of $l$ can detect more candidate images with poisonous triggers. (2) In the candidate trigger detection phase, the total processing time consistently increases and the average processing time performs relatively stable while the cluster count of $l$ increases. The system will spend more calculating time with a larger cluster count of $l$. To conclude, an appropriate value of $l$ should be chosen to balance the performance and time costs. We empirically set $l=1000$.
VI-H Analysis of Backdoored SSL Process
In our assessment of the backdoored SSL process on ImageNet-100 with a poison rate of $0.5\%$, we conduct a thorough investigation into the correlation between the overall model performance, as measured by Clean Data Acc, and the attack effectiveness, as measured by the Patched Data FP. As shown in Table III, we evaluate the attack effectiveness using poisoned data to train the self-supervised model with different training epochs. We test the accuracy and false positive rates of the model on clean and poisoned data after fine-tuning on downstream classification tasks using clean datasets. (poison rate 0.5%, target category “laptop” and “goose”). The results show a strong correlation between the effectiveness of the attack and the model performance before 100 epochs. Our findings indicate that in the initial epochs ($epoch\leq 40$), attack effectiveness has a strong correlation with overall model performance. Although this correlation is not as stringent in later epochs, the attack effectiveness still maintains its overall magnitude. These observations suggest that, when training a model for defense, achieving convergence may not be necessary. Instead, it is sufficient to train the model until the attack effectiveness is comparable to that of the fully trained model in overall magnitude.
VII Conclusion